OpenStudy (anonymous):

Fayed tipped the pizza delivery driver two dollars for pizzas that cost twenty-four dollars. Approximately what percent did he tip the driver? Round your answer to the nearest percent.

OpenStudy (anonymous):

Correct answer gets medal and follow

OpenStudy (anonymous):

8.3 %

OpenStudy (anonymous):

Are you sure? I really need to get a good grade on this assignment

OpenStudy (anonymous):

use proportions is 2/24=x/100 2*100=200 200/24=8.3

OpenStudy (anonymous):

Ok thanks!